# Polyp SAM 2: Advancing Zero shot Polyp Segmentation in Colorectal Cancer Detection

### v2

@@ 1 Introduction
The Segment Anything Model (SAM) [17] has demonstrated remarkable success in zero-shot image segmentation tasks. SAM’s prompt-based approach allows users to specify objects of interest without the need for additional training. By producing high-quality object masks from input prompts (such as points, boxes, and masks), SAM has garnered attention for various applications. For example, [26] assessed SAM’s zero-shot capabilities in organ segmentation tasks. Reference [31] specifically evaluated SAM’s performance in Camouflaged Object Detection (COD). Meanwhile, Reference [36] explored the benefits and limitations of the SAM in both computer vision and medical image segmentation tasks. Additionally, [46] evaluated the performance of SAM in segmenting polyps, in which SAM is under unprompted settings. Deng et al. [5] applied SAM to segment heterogeneous objects in digital pathology.
The Segment Anything Model 2 (SAM 2) [7] builds upon the success of its predecessor, SAM, addressing the challenges posed by real time image and video segmentation. Trained on video data, SAM 2 offers real-time capabilities, allowing it to segment entire video sequences based on annotations from a single frame. By leveraging interactions (such as clicks, boxes, or masks) across frames, SAM 2 predicts spatial-temporal masks (referred to as ‘masklets’) for objects. Its unified approach reduces user interaction time and enhances performance, making it a promising tool for various applications, including medical image segmentation and video analysis [36, 27, 49, 23].

### v4

@@ 1 Introduction
The Segment Anything Model (SAM) [17] has demonstrated remarkable success in zero-shot image segmentation tasks. SAM’s prompt-based approach allows users to specify objects of interest without the need for additional training. By producing high-quality object masks from input prompts (such as points, boxes, and masks), SAM has garnered attention for various applications. For example, [27] assessed SAM’s zero-shot capabilities in organ segmentation tasks. [32] specifically evaluated SAM’s performance in Camouflaged Object Detection (COD). Meanwhile, [15] explored the benefits and limitations of the Segment-Anything Model (SAM) in both computer vision and medical image segmentation tasks. Additionally, [47] evaluated the performance of SAM in segmenting polyps, in which SAM is under unprompted settings. Deng et al. [5] applied SAM to segment heterogeneous objects in digital pathology.
The Segment Anything Model 2 (SAM 2) [25] builds upon the success of its predecessor, SAM, addressing the challenges posed by real time image and video segmentations. Trained on video data, SAM 2 offers real-time capabilities, allowing it to segment entire video sequences based on annotations from a single frame. By leveraging interactions (such as clicks, boxes, or masks) across frames, SAM 2 predicts spatial-temporal masks (referred to as ‘masklets’) for objects. Its unified approach reduces user interaction time and enhances performance, making it a promising tool for various applications, including medical image segmentation and video analysis [37, 28, 50, 23].